# Clarification and Coinduction of Tambara Functors

### v1

@@ 5.3. Morita invariance and $K$-theory
$k\rightarrow\mathrm{Coind}_{H}^{G}\mathrm{Res}_{H}^{G}k\mathrm{.}$
@@
$f^{*}\oplus g^{*}:R{-}\mathsf{Mod}\oplus S{-}\mathsf{Mod}\rightarrow R\times S{-}\mathsf{Mod}$
@@
Since limits are computed levelwise, we have $(R\times S)(G/H)\cong R(G/H)\times S(G/H)$. One checks straighftorwardly using Proposition 2.21 that this induces an $R\times S$-module isomorphism $M\cong M^{\prime}\oplus M^{\prime\prime}$ for any $R\times S$-module $M$, where $M^{\prime}$ is in the image of restriction along $f$ and $M^{\prime\prime}$ is in the image of restriction along $g$. Therefore $f^{*}\oplus g^{*}$ is essentially surjective, and one sees that every morphism $f^{*}M\oplus g^{*}N\rightarrow f^{*}M^{\prime}\oplus g^{*}N^{\prime}$ is diagonal, i.e. $f^{*}\oplus g^{*}$ is full.
@@
$R{-}\mathsf{Mod}\rightarrow\mathrm{Coind}_{H}^{G}R{-}\mathsf{Mod}\mathrm{.}$
@@
$M(G/L)\cong\prod_{g\in H\backslash G/L}d_{g,L}M(G/L)\mathrm{.}$
@@
First, observe that $\phi$ becomes an isomorphism after applying $\mathrm{Res}_{H}^{G}$. The argument is essentially identical to the argument of Lemma 3.8, except that we replace the ring structure everywhere with the structure of a module over $\mathrm{Coind}_{H}^{G}R$; the point is that the conjugations in $\mathrm{Coind}_{H}^{G}R$ permute the idempotents which pick out the various factors of $M(G/L)$, hence the conjugations in $M$ determine the appropriate isomorphisms between the factors obtained after applying $\mathrm{Res}_{H}^{G}$. The argument of Proposition 3.9 also goes through, and shows that $\phi$ is itself an isomorphism; the main point is that injectivity is forced by Frobenius reciprocity in exact the same way, using the module structure rather than the ring structure. Finally, one checks that $\phi$ is an isomorphism of $\mathrm{Coind}_{H}^{G}R$-modules. Therefore $\mathrm{Coind}_{H}^{G}$ is essentially surjective.
@@
$\oplus_{i}R_{i}{-}\mathsf{Mod}\cong\oplus_{j}R_{j}{-}\mathsf{Mod}$
@@
Let $R$ be any field-like or domain-like $G$-Tambara functor. Then the $K$-theory of $R$ is isomorphic to the $K$-theory of, respectively, a field-like or domain-like clarified $H$-Tambara, for some $H\subset G$.
@@
This follows immediately from Corollary 3.10, Proposition 5.27, Proposition 5.7, and the fact that $K$-theory only depends on the category of modules (in particular, the exact category of finitely generated projective modules, which is preserved by Morita equivalence). ∎
In fact, Proposition 5.7 may be pushed further, to the category of algebras in Green or Tambara functors. Because Green functors are precisely commutative monoids over the box product, we immediately deduce an analogue of Proposition 5.7 for Green functors (namely, if $k$ is a Green functor, then the category of $k$-algebras is equivalent to the category of $\mathrm{Coind}_{H}^{G}k$-algebras, in Green functors). Before we can prove the analogous result for Tambara functors, we require a lemma.
@@
$H{-}\mathsf{Tamb}_{k/}\cong G{-}\mathsf{Tamb}_{\mathrm{Coind}_{H}^{G}k/}$
@@
By Proposition 5.7 and the fact that Green functors are precisely the commutative monoids with respect to the box product, we observe that coinduction induces an equivalence between the respective algebra categories of Green functors. By Proposition 5.11, $\mathrm{Coind}_{H}^{G}:H{-}\mathsf{Tamb}_{k/}\cong G{-}\mathsf{Tamb}_{\mathrm{Coind}_{H}^{G}k}$ is essentially surjective. Faithfulness follows from faithfulness in the Green functor setting and the fact that the forgetful functor from Tambara to Green functors is faithful.

### v2

@@ 5.1. Morita invariance and $K$-theory
$k\rightarrow\mathrm{Coind}_{H}^{G}\mathrm{Res}_{H}^{G}k.$
@@
$f^{*}\oplus g^{*}:R{-}\mathrm{Mod}\oplus S{-}\mathrm{Mod}\rightarrow R\times S{-}\mathrm{Mod}$
@@
Since limits are computed levelwise, we have $(R\times S)(G/H)\cong R(G/H)\times S(G/H)$. One checks straighftorwardly using the Dress pairing description of relative box products (cf. [CW25]) that this induces an $R\times S$-module isomorphism $M\cong M^{\prime}\oplus M^{\prime\prime}$ for any $R\times S$-module $M$, where $M^{\prime}$ is in the image of restriction along $f$ and $M^{\prime\prime}$ is in the image of restriction along $g$. Therefore $f^{*}\oplus g^{*}$ is essentially surjective, and one sees that every morphism $f^{*}M\oplus g^{*}N\rightarrow f^{*}M^{\prime}\oplus g^{*}N^{\prime}$ is diagonal, i.e. $f^{*}\oplus g^{*}$ is full.
@@
$R{-}\mathrm{Mod}\rightarrow\mathrm{Coind}_{H}^{G}R{-}\mathrm{Mod}.$
@@
$M(G/L)\cong\prod_{g\in H\backslash G/L}d_{g,L}M(G/L).$
@@
First, observe that $\phi$ becomes an isomorphism after applying $\mathrm{Res}_{H}^{G}$. The argument is essentially identical to the first half of the argument of Proposition 3.8, except that we replace the ring structure everywhere with the structure of a module over $\mathrm{Coind}_{H}^{G}R$; the point is that the conjugations in $\mathrm{Coind}_{H}^{G}R$ permute the idempotents which pick out the various factors of $M(G/L)$, hence the conjugations in $M$ determine the appropriate isomorphisms between the factors obtained after applying $\mathrm{Res}_{H}^{G}$. The argument of Proposition 3.8 also goes through, and shows that $\phi$ is itself an isomorphism; the main point is that injectivity is forced by Frobenius reciprocity in exact the same way, using the module structure rather than the ring structure. Finally, one checks that $\phi$ is an isomorphism of $\mathrm{Coind}_{H}^{G}R$-modules. Therefore $\mathrm{Coind}_{H}^{G}$ is essentially surjective.
@@
$\oplus_{i}R_{i}{-}\mathrm{Mod}\cong\oplus_{j}S_{j}{-}\mathrm{Mod}$
@@
Let $R$ be any field-like $G$-Tambara functor. Then the $K$-theory of $R$ is isomorphic to the $K$-theory of a clarified field-like $H$-Tambara, for some $H\subset G$.
@@
This follows immediately from Corollary 3.9, Proposition 5.4, and the fact that $K$-theory only depends on the category of modules (in particular, the exact category of finitely generated projective modules, which is preserved by Morita equivalence). ∎
In particular, note that Chan and the author in [CW25] study the algebraic $K$-theory of clarified Tambara functors (as a subclass of the class of Green meadows). Since every field-like Tambara functor is coinduced from a clarified field-like Tambara functor, the results in [CW25] actually apply to a much larger class of field-like Tambara functors.
Proposition 5.4 may be pushed further, to the category of algebras in Green or Tambara functors. Because Green functors are precisely commutative monoids over the box product, we immediately deduce an analogue of Proposition 5.4 for Green functors (namely, if $k$ is a Green functor, then the category of $k$-algebras is equivalent to the category of $\mathrm{Coind}_{H}^{G}k$-algebras, in Green functors). Before we can prove the analogous result for Tambara functors, we require a lemma.
@@
$H{-}\mathrm{Tamb}_{k/}\cong G{-}\mathrm{Tamb}_{\mathrm{Coind}_{H}^{G}k/}$
@@
By Proposition 5.4 and the fact that Green functors are precisely the commutative monoids with respect to the box product, we observe that coinduction induces an equivalence between the respective algebra categories of Green functors. By Proposition 5.8, $\mathrm{Coind}_{H}^{G}:H{-}\mathrm{Tamb}_{k/}\cong G{-}\mathrm{Tamb}_{\mathrm{Coind}_{H}^{G}k}$ is essentially surjective. Faithfulness follows from faithfulness in the Green functor setting and the fact that the forgetful functor from Tambara to Green functors is faithful.